2. Solving for the WACC

The WACC is used as the discount rate to evaluate various capital budgeting projects. However, it is important to realize that the WACC is an appropriate discount rate only for a project of average risk.

Analyze the cost of capital situations of the following company cases, and answer the specific questions that finance professionals need to address.

Consider the case of Turnbull Co.

Turnbull Co. has a target capital structure of 58% debt, 6% preferred stock, and 36% common equity. It has a before-tax cost of debt of 8.2%, and its cost of preferred stock is 9.3%.

If Turnbull can raise all of its equity capital from retained earnings, its cost of common equity will be 12.4%. However, if it is necessary to raise new common equity, it will carry a cost of 14.2%.

If its current tax rate is 25%, how much higher will Turnbull’s weighted average cost of capital (WACC) be if it has to raise additional common equity capital by issuing new common stock instead of raising the funds through retained earnings? (Note: Round your intermediate calculations to two decimal places.)

0.59%

0.65%

0.85%

0.88%

Turnbull Co. is considering a project that requires an initial investment of $270,000. The firm will raise the $270,000 in capital by issuing $100,000 of debt at a before-tax cost of 8.7%, $30,000 of preferred stock at a cost of 9.9%, and $140,000 of equity at a cost of 13.2%. The firm faces a tax rate of 25%.

What will be the WACC for this project? (Note: Round your intermediate calculations to three decimal places.)

A. 10.36%

B. 6.73%

C. 7.77%

D. 11.91%

Consider the case of Kuhn Co.

Kuhn Co. is considering a new project that will require an initial investment of $20 million. It has a target capital structure of 45% debt, 4% preferred stock, and 51% common equity. Kuhn has noncallable bonds outstanding that mature in 15 years with a face value of $1,000, an annual coupon rate of 11%, and a market price of $1555.38. The yield on the company’s current bonds is a good approximation of the yield on any new bonds that it issues. The company can sell shares of preferred stock that pay an annual dividend of $8 at a price of $92.25 per share.

Kuhn does not have any retained earnings available to finance this project, so the firm will have to issue new common stock to help fund it. Its common stock is currently selling for $33.35 per share, and it is expected to pay a dividend of $1.36 at the end of next year. Flotation costs will represent 3% of the funds raised by issuing new common stock. The company is projected to grow at a constant rate of 8.7%, and they face a tax rate of 25%. What will be the WACC for this project? ____________ (Note: Round your intermediate calculations to two decimal places.)

A. 8.78%

B. 7.02%

C. 7.90%

D. 10.54%
Answer #1
Case-1
WACC if it has to raise additional common equity capital by issuing new common stock
(Wt.d*kd)+(Wt.ps*k ps)+(Wt.e*ke)
(58%*8.2%*(1-25%))+(6%*9.3%)+(36%*14.2%)=
9.24%
WACC if it has to raise additional common equity capital by raising from retained earnings
(58%*8.2%*(1-25%))+(6%*9.3%)+(36%*12.4%)=
8.59%
Amt. by which the latter exceeds the former
(9.24%-8.59%)=
0.65%
ANSWER: 2.----- 0.65%
Case--2
WACC = (Wt.d*kd)+(Wt.ps*k ps)+(Wt.e*ke)
ie.(100000/270000*8.7%*(1-25%))+(30000/270000*9.9%)+(140000/270000*13.2%)=
10.36%
ANSWER: a. 10.36%
3.Kuhn Co.
After-tax cost of debt:
Using the PV formula for bond price,
1555.38=((1000*11%)*(1-(1+r)^-15)/r)+(1000/(1+r)^15)
Solving for r, we get the before -tax yield on the bond as
5.48%
So, the after-tax yield on the bond as
5.48%*(1-25%)=
4.11%
Cost of preferred stock:
Annual $ dividend/Current pref. stock price
8/92.25=
8.67%
Cost of new equity using Gordon's constant dividend growth model,
ke=(D1/P0)+g
where.
ke= cost of equity
D1= the next dividend,ie.D0*(1+g)
ie.D1=1.36
P0= the current price--here, current price - Flotation costs,ie. 33.35*(1-3%)= 32.3495
g= the constant growth rate, ie. 8.7% or 0.087
Now, the cost of equity, ke=
(1.36/32.3495)+0.087=
12.90%
Now the WACC =
(Wt.d*kd)+(Wt.ps*k ps)+(Wt.e*ke)
ie. (45%*4.11%)+(4%*8.67%)+(51%*12.90%)=
8.78%
ANSWER: A--- 8.78%
The WACC for this project= 8.78%